Trash:

2x2-13x+7=x2 Round to the nearest tenth

Trash:

1 attachment
SaltTheLoser:

choose one, graphing or quadratic formula

Trash:

quadratic formula

jhonyy9:

like a first step subtract from both sides x^2

SaltTheLoser:

\[\frac{ -b \pm \sqrt{b^2-4ac} }{ 2a }\\ax^2+bx+c\\x^2-13x\]

SaltTheLoser:

-13x+7 * plug it in, and do it

Trash:

Alr ty

jhonyy9:

@salttheloser step by step please and explain it

SaltTheLoser:

@trash do you get how to do it by yourself?

Trash:

Yezzir thank yall<3

surjithayer:

\[2x^2-13x+7=x^2\] or \[x^2-13x+7=0\] either solve by completing squares or by using quadratic formula \[x^2-13x+(-\frac{ 13 }{ 2 })^2=-7+(-\frac{ 13 }{ 2 })^2\] \[(x-\frac{ 13 }{ 2 })^2=-7+\frac{ 169 }{ 4 }=\frac{ -28+169 }{ 4 }=\frac{ 141 }{ 4 }\] \[x=\frac{ 13 }{2 }\pm \frac{ \sqrt{141} }{ 2 }\] \[either~x=\frac{ 13+\sqrt{141} }{ 2 }\approx 12.44\] \[or~x=\frac{ 13-\sqrt{141} }{ 2 }\approx 0.56\]